OpenStudy (gracygirl):

A "Local" train leaves a station and runs at an average rate of 35 mph. An hour and a half later an "Express" train leaves the station and travels at an average rate of 56 mph on a parallel track. How many hours after the Express train starts will the it overtake the Local?

OpenStudy (gracygirl):

@Will.H @AloneS

OpenStudy (frozen15):

what do you think? what equations or ideas did you use??

OpenStudy (gracygirl):

I dont know :/ not good at these

OpenStudy (frozen15):

okay Here's what I did local=L, express=E L at 35mph will travel 35(t) miles in t hours E at 56mph will travel 56(t) miles in t hours you wanna know when they are at the same distance, but E 1:30 hr later, this means the local train is already 35(1.5) miles ahead, 52.5 miles.. so, 35(t)+52.5 = 56(t) 52.5=56t-35t 52.5=21t t=52.5/21=2.5, two hrs and a half later they are at the same distance Makes sense??

OpenStudy (gracygirl):

Yeah, thank you!

OpenStudy (frozen15):

your welcome!! I gtg so I can no longer help , good luck on future questions!